OpenStudy (anonymous):

Linear Algebra question: For any matrix A, why does rank(A)=rank(A^T)? Any good proofs?

OpenStudy (anonymous):

here is a short one col rank (A) = row rank (A^T) so dim (A) = col rank (A) = row rank (A^T) = rank (A^T)

OpenStudy (anonymous):

What exactly is dim(A)? Dimension of the column space?